A bowling ball with a mass of 5.4 kg moves in a straight line at 2.4 m/s. While moving in a straight line, how fast must a 0.001143 kg ping-pong ball travel so that the two end up with the same momentum ?


1
Expert's answer
2021-01-28T19:38:53-0500
"p_1=p_2,""m_1v_1=m_2v_2,""v_2=\dfrac{m_1v_1}{m_2}=\dfrac{5.4\ kg\cdot 2.4\ \dfrac{m}{s}}{0.001143\ kg}=11338.58\ \dfrac{m}{s}."